# STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

### v1

@@ Abstract
A simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) enhanced unnamed aerial vehicle (UAV)-enabled multi-user mobile edge computing (MEC) scheme is proposed in this paper. Different from the existing MEC works, the proposed scheme allows bi-directional offloading where users can simultaneously offload their computing tasks to the MEC servers situated at the ground base station (BS) and aerial UAV with the assistance of the STAR-RIS. Specifically, we formulate an optimization problem aiming at maximizing the energy efficiency of the system while ensuring the quality of service (QoS) constraint by jointly optimizing the resource allocation, user scheduling, passive beamforming of the STAR-RIS, and the UAV trajectory. An iterative algorithm designed with the Dinkelbach’s algorithm and the successive convex approximation (SCA) is proposed to effectively handle the formulated non-convex optimization problem. Simulation results indicate that the proposed STAR-RIS enhanced UAV-enabled MEC scheme possesses significant advantages in enhancing the system energy efficiency over other baseline schemes including the conventional RIS-aided scheme.
@@
STAR-RIS, unnamed aerial vehicle (UAV), mobile edge computing (MEC), energy efficiency.
@@ I Introduction
Nowadays, the number of wireless devices is experiencing an exponential explosive growth, and the applications of the Internet of Things (IoT) are becoming increasingly diverse. As for the implementation of those computation-intensive and latency-critical applications, e.g., autonomous driving, augmented and virtual reality, etc., it presents significant challenges for the widely used center-cloud computing framework to effectively handle large amounts of data in swift actions. [1, 2, 3]. To address this challenge, the technology of mobile edge computing (MEC) has emerged as a promising solution for effectively tackling the significantly augmented computational necessity, which is able to bring the capabilities of cloud computing to the network edge. The technology of MEC enables data processing in close proximity, and thus it performs well in reducing network congestion, and improving service quality and user experience [4, 5, 6]. While the MEC technology provides an effective means to enhance the network computing capabilities, the traditional placement strategy for MEC servers near the ground base stations (BSs) or access points (APs) may result in a limited service coverage.
@@
To overcome this limitation, the unmanned aerial vehicles (UAVs) are leveraged to assist the task completion of the MEC networks due to the inherent advantages of UAVs, such as exceptional mobility and flexibility [7, 8, 9, 10, 11, 12, 13, 14]. Specifically, in [7], the UAV is equipped with a MEC server acting as an aerial MEC platform to facilitate the computation of the offloaded tasks for users with low-quality transmission links from the BSs or APs. In [8], the UAV is leveraged as a relay to support the transfer of users’ computational tasks to the MEC servers located at BS. The authors in [9] introduce a novel MEC scheme that involves both aerial and ground cooperation. The proposed scheme enables users to efficiently offload their task data to multiple base stations (BSs) and unmanned aerial vehicles (UAVs) in a collaborative manner. Furthermore, a two-way offloading UAV-aided MEC scheme is proposed in [10, 12, 11] to enhance the computing capacity of the MEC network, where the UAV not only carries the MEC processor but also serves as a relay to facilitate the offloading of tasks from users to ground MEC servers. Although UAV is able to effectively improve the computing capacity of MEC networks, the existing UAV-aided MEC schemes are designed by adapting to uncontrollable random wireless channels, which seriously limits the task offloading efficiency.
@@
A promising solution to tackle this issue is the reconfigurable intelligent surface (RIS) technique [15]. Due to the fact that RIS can dynamically adjust the phases and amplitudes of incident signals, allowing the creation of controllable end-to-end virtual channels, RIS has been incorporated into various kinds of communication systems including MEC networks [16, 17, 18, 19]. It is important to highlight that UAVs flying at high altitudes enables the establishment of a reliable line-of-sight (LoS) connection between the UAV and users with a high possibility. Additionally, RIS technology has the capability to reconfigure the wireless propagation environment, and thus combining UAV and RIS will be a win-win strategy for MEC networks. In particular, a RIS-assisted UAV-enabled MEC scheme considering the aim of maximizing energy efficiency is proposed in [17]. In [18], the UAV-mounted RIS scheme is proposed to assist MEC network, where users can only offload their tasks to ground MEC servers with the help of the reflected ability of the RIS. To further improve the computational capacity of the MEC network, a two-way offloading scheme assisted by the UAV and the RIS is proposed in [19, 20]. In this scheme, a multi-antenna UAV is responsible for processing a portion of the computation tasks and acts as a relay to transmit the remaining tasks to the BS in the next time slot.
Actually, the existing UAV-enabled MEC schemes usually leverage a two-step task offloading strategy to achieve two-way task offloading to the BS and UAV, which have certain limitations: (i) The UAV serves as multiple roles where it acts as an MEC platform for processing partial offloaded tasks, and a relay to send the remaining tasks to the BS, which imposes a challenge on UAV hardware design. (ii) These two-step MEC schemes are not energy and time efficient since the UAV needs to receive all the offloaded tasks, and then transmit the unprocessed tasks to the BS. Hence, it is challenging for the traditional RIS-assisted UAV-enabled MEC schemes to offload users’ tasks to both the BS and the UAV simultaneously in a more efficient manner [19, 20]. Recently, an advanced RIS technology, named as simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS), has drawn great attention from both academia and industry[21, 22]. STAR-RIS can split the incident signal into two parts, where one part is reflected to the same side of the incident signal and the other part is transited to the opposite side, allowing a ${}^{\circ}$ coverage, compared with the conventional RIS. Due to the inherent advantages, STAR-RIS possesses an enormous application potentials in various wireless communication systems, e.g.,[23, 24, 25, 26]. Hence, to enhance the current MEC schemes, we try to incorporate the STAR-RIS into UAV-assisted MEC networks, allowing the simultaneously bi-directional offloading strategy for the cooperation between the UAV and the BS. In this paper, our main contributions are summarized as follows:
@@
  • STAR-RIS enhanced UAV-enabled MEC Architecture with Bi-directional Offloading: In this paper, a novel MEC scheme aided by STAR-RIS and UAV is proposed for the first time. In contrast to the existing UAV-enabled MEC schemes, the proposed scheme allows users to simultaneously offload their computing tasks to the MEC servers located at the BS and UAV in a bi-directional manner through the reflection and transmission capabilities of the STAR-RIS. Note that the UAV in the proposed MEC scheme is solely responsible for carrying the MEC server and the STAR-RIS to assist the computation and offloading of users’ tasks.
@@
  • Optimization Problem Formulation Maximizing Energy Efficiency under Practical Constraints: We formulate optimization problem with the aim of maximizing the energy efficiency of the system and ensuring users’ quality of service (QoS) constraints by jointly designing the resource allocation, users scheduling, passive beamforming and UAV trajectory. Actually, managing this optimization problem can be quite difficult because of the presence of a fractional objective function and the significant couplings among optimization variables.
@@
  • Iterative Algorithm with Guaranteed Convergence and Substantial Performance Gain: To effectively address this non-convex optimization problem, the alternative strategy is leveraged to divide the optimization problem into three subproblems. Then, an iterative algorithm based on the Dinkelbach’s algorithm [27, Chapter 3.2.1] and the successive convex approximation (SCA) technique is proposed to effectively solve these three subproblems. The convergence of the proposed algorithm can be guaranteed, which has also been confirmed through simulation results. Moreover, the performance improvement and potential applications of the STAR-RIS enhanced UAV-enabled MEC scheme are demonstrated by comparing the simulation results with three other baseline schemes, including the conventional RIS-aided scheme.
@@ II System Model
Fig. 1 shows the UAV-enabled MEC network assisted by STAR-RIS, which mainly consists of a ground base station (BS), $K$ ground users each with a single antenna, a aerial UAV equipped with a signal antenna and mounted a STAR-RIS with $M$ elements, two MEC servers respectively located at the BS and UAV. In this letter, the energy splitting protocol is executed for STAR-RIS, signifying that all elements integrated into the STAR-RIS possess the capability to simultaneously reflect (R) and transmit (T) the incident signals [21]. This characteristic implies that users can offload their computing tasks in a bi-directional manner, simultaneously to the MEC server located at the BS using the R mode of the STAR-RIS, as well as to the MEC server mounted on the UAV using the T mode.
@@
In this paper, we divide the mission period $T$ into $N$ equal time slots, i.e., $\delta_{\mathrm{t}}=T/N$, which is sufficiently small. We assume that the UAV is equipped with a transceiver that has a single antenna and the time division multiple access (TDMA) protocol is implemented to handle users’ offloaded tasks, indicating that only one user will be chosen to offload their computing tasks to the BS and the UAV within a time slot. Hence, we use the variable $\zeta_{k}[n]\in\{0,1\}$ for $k\in\mathcal{K}\triangleq\{1,\cdots,k,\cdots,K\},n\in\mathcal{N}\triangleq\{1,\cdots,n,\cdots,N\}$ to represent the user association decision for task offloading in a time slot. In particular, if $\zeta_{k}[n]=1$, it means that the $k$-th user is chosen to offload its task to the BS and UAV in the $n$-th time slot with the assistance of the STAR-RIS. Additionally, variable $\zeta_{k}[n]$ needs to satisfy the following constraint:
$\sum\limits_{k=1}^{K}\zeta_{k}[n]=1,~{}\zeta_{k}[n]\in\{0,1\},~{}\forall k\in\mathcal{K},n\in\mathcal{N}.$ (1)
@@
$\mathbf{v}_{\mathrm{ua}}[n]=\frac{\mathbf{q}_{\mathrm{ua}}[n+1]-\mathbf{q}_{\mathrm{ua}}[n]}{\delta_{\mathrm{t}}},\left\|\mathbf{v}_{\mathrm{ua}}[n]\right\|\leq v_{\max},$ (2)
$\mathbf{a}_{\mathrm{ua}}[n]=\frac{\mathbf{v}_{\mathrm{ua}}[n+1]-\mathbf{v}_{\mathrm{ua}}[n]}{\delta_{\mathrm{t}}},\left\|\mathbf{a}_{\mathrm{ua}}[n]\right\|\leq a_{\max},$ (3)
$\mathbf{q}_{\mathrm{ua}}[0]=\mathbf{q}_{\mathrm{I}},\mathbf{q}_{\mathrm{ua}}[N+1]=\mathbf{q}_{\mathrm{F}},$ (4)
where $\left\|\mathbf{v}_{\mathrm{ua}}[n]\right\|$ and $\left\|\mathbf{a}_{\mathrm{ua}}[n]\right\|$ respectively represent the flight speed and the acceleration of the UAV in the $n$-th time slot, with the maximum values of $v_{\max}$ and $a_{\max}$. In terms of $\mathbf{q}_{\mathrm{I}}$ and $\mathbf{q}_{\mathrm{F}}$, these locations can serve as the ground station where the UAV can access a reliable power supply and receive necessary maintenance.
@@ II-A Channel Model
$\mathbf{h}_{\varsigma}[n]=\sqrt{\frac{\rho}{d_{\varsigma}^{\alpha_{\varsigma}}[n]}}\widehat{\mathbf{h}}_{\varsigma}[n],~{}\varsigma\in\{\mathrm{r}k,\mathrm{rb}\},$ (5)
@@
with $\rho=(\frac{\lambda}{4\pi})^{2}$ and $\lambda$ being the wavelength of the carrier frequency, $\alpha_{\varsigma}$ indicating the pass loss exponent, $d_{\varsigma}[n]$ denoting the distance between $k$-th user/BS and the STAR-RIS. $d$ denotes the adjacent element separation of the STAR-RIS. In addition, the $\xi_{\varsigma}[n]$ and $\chi_{\varsigma}[n]$ corresponding to the $k$-th user and the BS are respectively calculated as
@@
$\mathbf{h}_{\mathrm{ru}}=\boldsymbol{\alpha}\circ\mathbf{a}_{\mathrm{ru}},$ (11)
@@
with $r_{m}$ being the distance between the antenna at UAV and the $m$-th element equipped at STAR-RIS. It is worth noting that the channel $\mathbf{h}_{\mathrm{ru}}$ remains constant due to the fixed relative position between the UAV and the STAR-RIS.

### v3

# Energy-Efficient STAR-RIS Enhanced UAV-Enabled MEC Networks with Bi-Directional Task Offloading

@@ Abstract
This paper introduces a novel multi-user mobile edge computing (MEC) scheme facilitated by the simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS) and the unmanned aerial vehicle (UAV). Unlike existing MEC approaches, the proposed scheme enables bi-directional offloading, allowing users to concurrently offload tasks to the MEC servers located at the ground base station (BS) and UAV with STAR-RIS support. To evaluate the effectiveness of the proposed MEC scheme, we first formulate an optimization problem aiming at maximizing the energy efficiency of the system while ensuring the quality of service (QoS) constraints by jointly optimizing the resource allocation, user scheduling, passive beamforming of the STAR-RIS, and the UAV trajectory. A block coordinate descent (BCD) iterative algorithm designed with the Dinkelbach’s algorithm and the successive convex approximation (SCA) technique is proposed to effectively handle the formulated non-convex optimization problem with significant coupling among variables. Simulation results indicate that the proposed STAR-RIS enhanced UAV-enabled MEC scheme possesses significant advantages in enhancing the system energy efficiency over other baseline schemes including the conventional RIS-aided scheme.
@@
STAR-RIS, unmanned aerial vehicle (UAV), mobile edge computing (MEC), energy efficiency.
@@ I Introduction
Currently, there is an exponential surge in the proliferation of wireless devices, accompanied by a notable diversification in the utilization of Internet of Things (IoT) applications. As for the implementation of those computation-intensive and latency-critical applications, e.g., autonomous driving, augmented and virtual reality, etc., it presents significant challenges for the widely used center-cloud computing framework to effectively handle large amounts of data in swift actions [1, 2, 3]. To address this challenge, the technology of mobile edge computing (MEC) has emerged as a promising solution for effectively tackling the significantly augmented computational necessity, since it is able to bring the capabilities of cloud computing to the network edge and enables data processing in close proximity. Consequently the MEC technique can perform well in reducing network congestion, and improving service quality and user experience. The benefits introduced by MEC have garnered significant attention from researchers, and thus many efforts primarily focus on reducing latency, conserving energy, enhancing energy efficiency and so on [4, 5, 6, 7, 8].
I-A Related Works
@@
While the MEC technology provides an effective means to enhance the network computing capabilities, the traditional placement strategy for MEC servers near the ground base stations (BSs) or access points (APs) may result in a limited service coverage. To overcome this limitation, the unmanned aerial vehicles (UAVs) are leveraged to assist the task completion of the MEC networks due to the inherent advantages of UAVs, such as exceptional mobility and flexibility [9, 10, 11, 12, 13, 14, 15, 16, 17]. Specifically, in [9], the UAV is equipped with a MEC server acting as an aerial MEC platform to facilitate the computation of the offloaded tasks for users with low-quality transmission links from the BSs or APs. In [10], the UAV is leveraged as a relay to support the transfer of users’ computational tasks to the MEC servers located at BS. The authors in [11] introduce a novel MEC scheme that involves both aerial and ground cooperation. The proposed scheme enables users to efficiently offload their task data to multiple base stations (BSs) and UAVs in a collaborative manner. Furthermore, a two-way offloading UAV-aided MEC scheme is proposed in [12, 14, 13] to enhance the computing capacity of the MEC network, where the UAV not only carries the MEC processor but also serves as a relay to facilitate the offloading of tasks from users to ground MEC servers. Although UAV is able to effectively improve the computing capacity of MEC networks, the existing UAV-aided MEC schemes are designed by adapting to uncontrollable random wireless channels, which seriously limits the task offloading efficiency.
@@
The reconfigurable intelligent surface (RIS) technique emerges as a promising solution to address this challenge [18, 19, 20]. Due to the fact that RIS can dynamically adjust the phases and amplitudes of incident signals, allowing the creation of controllable end-to-end virtual channels, RIS has been incorporated into various kinds of communication systems including MEC networks [21, 22, 23, 24, 25, 26, 27, 28]. It is important to highlight that UAVs flying at high altitudes enables the establishment of a reliable line-of-sight (LoS) connection between the UAV and users with a high possibility. Additionally, RIS technology has the capability to reconfigure the wireless propagation environment, and thus combining UAV and RIS will be a win-win strategy for MEC networks. In particular, a RIS-assisted UAV-enabled MEC scheme considering the aim of maximizing energy efficiency is proposed in [24]. This MEC scheme utilizes the RIS positioned on the building to enhance the signals of users’ task offloading and direct them towards the MEC server located on the UAV. To further improve the computational capacity of the MEC network, a two-way offloading scheme assisted by the UAV and the RIS is proposed in [25]. In this scheme, a multi-antenna UAV is responsible for processing a portion of users’ computation tasks while also acts as a relay to transmit the remaining computation tasks to the BS with the assistance of the RIS installed on the building’s surface.
It is worth nothing that the flexibility and mobility of the RIS in traditional RIS-assisted UAV-enabled MEC schemes are limited [24, 25], considering that the location of the RIS is usually fixed. In order to improve the flexibility and mobility of the RIS, an aerial RIS-aided MEC scheme is proposed to assist the MEC network in [26]. In this scheme, users only offload their tasks to ground MEC server with the help of the reflected ability of the RIS. Actually, the utilization of computing resources on the UAV is significantly limited by the reflection-only RIS, as it redirects all signals intended for task offloading to the BS. A MEC scheme utilizing two UAVs has been introduced by Duo et al. in [27] to overcome this constraint. In this scheme, one UAV is equipped with a RIS to support and optimize the transmission of task offloading signals from users to the MEC server positioned on the second UAV. It is important to highlight that the UAV housing the MEC server is capable to handle some of the computational tasks and serve as a relay to forward the remaining tasks to the ground BS, thereby significantly enhancing the computational capacity of the MEC network.
Actually, in the existing RIS-aided wireless communication schemes, the traditional RIS is solely able to execute the reflection modulation to the incident signals, which requires the transceiver terminal equipment to be located in the same side of the RIS. In other words, the conventional RIS can only reconfigure the half-space wireless propagation environment, which will significantly limit the coverage area of wireless networks and the flexibility in deploying the RIS. To breakthrough this limitation, an advanced RIS technology, named as simultaneously transmitting and reflecting reconfigurable intelligent surface (STAR-RIS), has drawn great attention from both academia and industry[29, 30, 31, 32]. STAR-RIS can split the incident signal into two parts, where one part is reflected to the same side of the incident signal and the other part is transited to the opposite side, allowing a 360 coverage, compared with the conventional RIS. As shown in Fig. 1, we present the main difference between the STAR-RIS and RIS in signal processing. Specifically, when the BS transmits signals to the traditional RIS, the signals can only be reflected towards the user positioned on the same side as the BS. The user on the opposite side will not receive any signals from the RIS. However, the STAR-RIS has the capability to not only reflect signals to users positioned on the same side as the BS but also transmit signals to users on the opposing side.
Figure 1: The main difference between the STAR-RIS and the conventional RIS in signal processing.
I-B Motivation and Contributions
Due to the inherent advantages, STAR-RIS possesses an enormous application potentials in various wireless communication systems, e.g., secure communications [33, 34, 35], integrated sensing and communications [36, 37] and MEC networks [38, 39, 40]. Specifically, a novel STAR-RIS-assisted MEC scheme is proposed in [40], where the STAR-RIS is attached vertically on the UAV. With the assistance of STAR-RIS, the ground users distributed in a 360 manner can efficiently offload their tasks to the BS. It is worth noting that the traditional RIS mounted horizontally on the UAV can offer a 360 coverage for ground users, as demonstrated in [26]. However, the aerial STAR-RIS-aided MEC system provides greater flexibility in modulating incident signals through the use of transmission and reflection beamforming. Note that, the MEC scheme supported by the STAR-RIS in [40] still exhibits certain constraints: (i) The vertically positioned STAR-RIS will experience air resistance when the UAV is swiftly moving in the air. And the resistance becomes more pronounced as the surface area of the RIS increases, which will affect the flexibility of the UAV. (ii) The MEC scheme underutilizes the complete spatial modulation potential inherent in the STAR-RIS and it is difficult for this scheme to use the computing resources on the UAV.
Although the existing RIS-assisted UAV-enabled MEC schemes with the two-way task offloading can fully leverage the computing resources situated at the BS and UAV, they have the following deficiencies: (i) The UAV serves as multiple roles where it acts as an MEC platform for processing partial offloaded tasks, and a relay to send the remaining tasks to the BS, which imposes a challenge on UAV hardware design. (ii) These two-step MEC schemes are not energy and time efficient since the UAV needs to receive and decode all the offloaded tasks, and then transmit the unprocessed tasks to the BS in the next time slot. The primary driving force behind this study stems from the imperative need to tackle the aforementioned challenges. To overcome limitations of the current MEC schemes, we propose the STAR-RIS enhanced UAV-assisted MEC scheme. In this scheme, the STAR-RIS is attached on the UAV parallel to the ground, which allows users to offload their computing tasks to MEC servers located at the BS and UAV simultaneously, facilitated by the reflection and transmission features of the STAR-RIS. In this paper, our main contributions are summarized as follows:
@@ I-B Motivation and Contributions
  • STAR-RIS enhanced UAV-enabled MEC Architecture with Bi-directional Offloading: A novel MEC scheme aided by the STAR-RIS horizontally mounted on the UAV is proposed for the first time. In contrast to the existing MEC schemes, the proposed scheme allows users to simultaneously offload their computing tasks to the MEC servers located at the BS and UAV in a bi-directional manner through the reflection and transmission capabilities of the STAR-RIS. Note that the UAV in the proposed MEC scheme is solely responsible for carrying the MEC server and the STAR-RIS to assist the computation and offloading of users’ tasks.
@@
  • Optimization Problem Formulation Maximizing Energy Efficiency under Practical Constraints: To assess the effectiveness of the proposed MEC scheme, We first formulate an optimization problem with the aim of maximizing the energy efficiency of the system and ensuring users’ quality of service (QoS) constraints by jointly designing the resource allocation, users scheduling, passive beamforming and UAV trajectory. Actually, managing this optimization problem can be quite difficult because of the presence of a fractional objective function and the significant couplings among optimization variables.
@@
  • Iterative Algorithm with Guaranteed Convergence and Substantial Performance Gain: To effectively address this non-convex optimization problem, the alternative strategy is leveraged to divide the optimization problem into three subproblems. Then, an iterative algorithm based on the Dinkelbach’s algorithm [41, Chapter 3.2.1] and the successive convex approximation (SCA) technique is proposed to effectively solve these three subproblems. The assured convergence and efficacy of the algorithm under consideration can be validated through an analysis of its convergence curve and a comparative assessment against the semidefinite relaxation (SDR) technique. Moreover, the potential of the STAR-RIS enhanced UAV-enabled MEC scheme are demonstrated by comparing the simulation results with four other baseline schemes.
@@ II System Model
Fig. 2 illustrates the architecture of the UAV-enabled MEC network with the assistance of the STAR-RIS. The network comprises a ground base station (BS), $K$ users each equipped with a single antenna, a UAV equipped with a signal antenna and installed with a STAR-RIS featuring $M$ elements, along with two MEC servers situated respectively at the BS and the UAV. This study adopts the energy splitting protocol for the STAR-RIS, wherein all elements incorporated within the STAR-RIS possess the capability to simultaneously reflect (R) and transmit (T) incident signals [29]. Specifically, when signals from users are incident on STAR-RIS, a portion of the signals are reflected to the BS by STAR-RIS, while the remaining portion of the signals reach the UAV via the STAR-RIS. This feature enables users to offload their computing tasks in a bidirectional manner concurrently to the MEC servers located at the BS and the UAV, respectively.
@@
In this paper, we divide the mission period $T$ into $N$ equal time slots, i.e., $\delta_{\mathrm{t}}=T/N$, which is sufficiently small. Considering that the UAV is equipped with a transceiver that has a single antenna, the time division multiple access (TDMA) protocol is implemented to handle users’ offloaded tasks, requiring that only one user will be chosen to offload the computing tasks to the BS and the UAV within a time slot. Here, we use the variable $\zeta_{k}[n]\in\{0,1\}$ for $k\in\mathcal{K}\triangleq\{1,\cdots,K\},n\in\mathcal{N}\triangleq\{1,\cdots,N\}$ to represent the user association decision for task offloading in a time slot. In particular, if $\zeta_{k}[n]=1$, it means that the $k$-th user is chosen to offload its task to the BS and UAV in the $n$-th time slot with the assistance of the STAR-RIS. To ensure that only one user is selected to offload its tasks in each time slot, variable $\zeta_{k}[n]$ needs to satisfy the following constraint:
$\sum\limits_{k=1}^{K}\zeta_{k}[n]=1,~{}\zeta_{k}[n]\in\{0,1\},~{}\forall k\in\mathcal{K},~{}n\in\mathcal{N}.$ (1)
@@
$\mathbf{v}_{\mathrm{ua}}[n]=\frac{\mathbf{q}_{\mathrm{ua}}[n+1]-\mathbf{q}_{\mathrm{ua}}[n]}{\delta_{\mathrm{t}}},~{}\left\|\mathbf{v}_{\mathrm{ua}}[n]\right\|\leq v_{\max},$ (2)
$\mathbf{a}_{\mathrm{ua}}[n]=\frac{\mathbf{v}_{\mathrm{ua}}[n+1]-\mathbf{v}_{\mathrm{ua}}[n]}{\delta_{\mathrm{t}}},~{}\left\|\mathbf{a}_{\mathrm{ua}}[n]\right\|\leq a_{\max},$ (3)
$\mathbf{q}_{\mathrm{ua}}[0]=\mathbf{q}_{\mathrm{I}},~{}\mathbf{q}_{\mathrm{ua}}[N+1]=\mathbf{q}_{\mathrm{F}},$ (4)
where $\left\|\mathbf{v}_{\mathrm{ua}}[n]\right\|$ and $\left\|\mathbf{a}_{\mathrm{ua}}[n]\right\|$ respectively represent the flight speed and the acceleration of the UAV in the $n$-th time slot, with the maximum values of $v_{\max}$ and $a_{\max}$. In terms of $\mathbf{q}_{\mathrm{I}}$ and $\mathbf{q}_{\mathrm{F}}$, they can serve as the ground station where the UAV can access a reliable power supply and receive necessary maintenance.
@@ II-A Channel Model
$\mathbf{h}_{\varsigma}[n]=\sqrt{\frac{\rho}{d_{\varsigma}^{\alpha_{\varsigma}}[n]}}\widehat{\mathbf{h}}_{\varsigma}[n]\in\mathbb{C}^{M\times 1},~{}\varsigma\in\{\mathrm{r}k,\mathrm{rb}\},$ (5)
@@
with $\rho=(\frac{\lambda}{4\pi})^{2}$ [43] represents the path loss at a reference distance of $1$ meter (m) with $\lambda$ being the wavelength of the carrier frequency, $\alpha_{\varsigma}$ indicating the pass loss exponent, $d_{\varsigma}[n]$ denoting the distance between $k$-th user/BS and the STAR-RIS. $d$ denotes the adjacent element separation of the STAR-RIS. In addition, the $\xi_{\varsigma}[n]$ and $\chi_{\varsigma}[n]$ corresponding to the $k$-th user and the BS are respectively calculated as
@@
$\mathbf{h}_{\mathrm{ru}}=\boldsymbol{\alpha}\circ\mathbf{a}_{\mathrm{ru}}\in\mathbb{C}^{M\times 1},$ (11)
@@
with $r_{m}$ denotes the distance between the antenna located on the UAV and the $m$-th element of the STAR-RIS, where $m\in\mathcal{M}\triangleq\{1,\cdots,M\}$. Notably, the channel $\mathbf{h}_{\mathrm{ru}}$ remains invariant throughout, attributed to the fixed spatial relationship maintained between the UAV and the STAR-RIS.